### v4

# AlphaDPO: Adaptive Reward Margin for Direct Preference Optimization

Abstract
Aligning large language models (LLMs) with human preferences requires balancing policy optimization with computational stability. While recent offline methods like DPO and SimPO bypass reinforcement learning’s complexity, they face critical limitations: DPO relies on static reference models that degrade with policy updates, and SimPO assumes a uniform target reward margin that ignores instance-wise preference strength. We propose AlphaDPO, an adaptive preference optimization framework that dynamically reparameterizes the reference distribution to address these issues. Our key innovation lies in an implicit reference model $\hat{\pi}_{\text{ref}}\propto U(y|x)(\pi_{\theta}/\pi_{\text{ref}})^{\alpha}$, which interpolates between policy-driven specialization and uniform exploration while enabling instance-adaptive reward margins. Theoretically, we prove AlphaDPO implicitly controls sequential KL divergence between iterative policy updates, ensuring stability even with poorly calibrated reference models. Empirically, AlphaDPO achieves state-of-the-art performance on AlpacaEval 2 (58.7% LC win rate) and Arena-Hard (35.7% win rate) across Mistral2-7B, Llama3-8B, and Gemma2-9B, demonstrating robust alignment without multi-stage training. Our work establishes adaptive reference reparameterization as a principled mechanism for preference optimization. The code is available at https://github.com/junkangwu/alpha-DPO.
Machine Learning, ICML
1 Preliminaries
Offline Alignment. In the offline alignment problem, we have access to a dataset $\mathcal{D}=\{(x,y_{w},y_{l})\}$ comprising prompts $x$ and labeled response pairs $(y_{w},y_{l})$ obtained from a reference policy $\pi_{\text{ref}}$. Here, $y_{w}$ is the preferred (winning) response and $y_{l}$ is the less preferred (losing) response. Although the underlying latent reward function $r^{*}(x,y)$ that governs these preferences is not directly observable, the Bradley-Terry (BT) model (Bradley & Terry, 1952) provides a framework for modeling pairwise comparisons:
$\mathbb{P}(y_{w}\succ y_{l}|x)=\frac{\exp(r^{*}(x,y_{w}))}{\exp(r^{*}(x,y_{w}))+\exp(r^{*}(x,y_{l}))},$ (1)
where $r^{*}(x,y)$ assigns a latent reward to each response $y$ given prompt $x$. The goal of offline alignment is to learn a policy $\pi_{\theta}$ that approximates $r^{*}(x,y)$ using $\mathcal{D}$.
Reinforcement Learning from Human Feedback (RLHF). Classical offline alignment algorithms employ reinforcement learning with a KL-regularized reward objective (Bai et al., 2022; Ziegler et al., 2019; Ouyang et al., 2022), defined for a regularization parameter $\eta>0$:
$\mathop{\max}_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}[r_{\phi}(x,y)]-\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)],$ (2)
where $r_{\phi}(x,y)$ is the reward function learned using the BT model on the preference dataset, $\pi_{\theta}$ is the policy model being optimized, $\pi_{\text{ref}}$ is the fixed reference policy, typically obtained via supervised fine-tuning. The KL-divergence regularizes the policy to remain close to the reference model.
Directed Preference Optimization (DPO). DPO (Rafailov et al., 2023) is a leading offline preference optimization method. Instead of learning an explicit reward model, DPO reparameterizes the reward function $r(x,y)$ using a closed-form expression involving the optimal policy:
$r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}+\beta\log Z(x),$ (3)
where $Z(x)$ is the partition function independent of $y$. This leads to the DPO loss for any triplet $(x,y_{w},y_{l})$:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (4)
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right].$
where $\sigma(\cdot)$ denotes the sigmoid function.
Simple Preference Optimization (SimPO). SimPO (Meng et al., 2024) introduces two key contributions: (1) a length-normalized reward, calculated as the average log-probability per token of a response under the policy model $\pi_{\theta}$, and (2) a target reward margin $\gamma$ to ensure the reward difference between winning and losing responses exceeds this margin. The SimPO loss is formulated as:
${\mathcal{L}}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (5)
$\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$
where $|y|$ denotes the length (i.e., number of tokens) of response $y$, normalizing the reward by response lengths, and $\gamma$ is the target reward margin.
2 Method
In this section, we establish a unified framework that connects DPO and SimPO (Section 2.1), highlighting the critical role of the reference model in preference optimization. We then introduce AlphaDPO (Section 2.2), a new preference optimization algorithm that synergizes the strengths of both DPO and SimPO.
2.1 A Common Framework for DPO and SimPO
A key insight in our work is that SimPO implicitly adopts a uniform distribution over the vocabulary as its reference model, whereas DPO employs the SFT model as the reference. By examining the role of the reference model in both methods, we derive the following result:
Theorem 2.1.
Let $U(y|x)$ denote a uniform distribution over the vocabulary for a given input $x$, replacing $\pi_{\text{ref}}(y|x)$ in the DPO loss function. Then, the DPO loss function simplifies to:
$\mathcal{L}(\pi_{\theta};U)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (6)
$\left[\log\sigma\left(\beta\left(\log\pi_{\theta}(y_{w}|x)-\log\pi_{\theta}(y_{l}|x)\right)-\gamma\right)\right],$
where $\gamma=\beta\left(\log U(y_{w}|x)-\log U(y_{l}|x)\right)$ is a constant. Under a length-normalized reward formulation, this loss function becomes:
$\mathcal{L}_{\text{LN}}(\pi_{\theta};U)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (7)
$\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right].$
Therefore, SimPO can be interpreted as a special case of DPO where the reference model is a uniform distribution.
Remark 2.2.
Why do winning and losing responses have different probabilities under a uniform policy? Although both winning and losing responses are sampled from the same policy model, their selection probabilities diverge through the reward model’s scoring mechanism. Specifically, winning response ($y_{w}$) are those assigned higher scores, while losing response ($y_{l}$) receive lower scores. Even under a uniform policy, these selections remain distinct because their selection probabilities are tied to their relative scores, not to the underlying uniform distribution over the response space.
Theorem 2.1 establishes a unified framework for DPO and SimPO by showing that replacing the reference model $\pi_{\text{ref}}$ in DPO with a uniform distribution $U$ reduces the DPO loss to the SimPO loss, up to a constant term $\gamma$. This reveals that SimPO is essentially DPO with a uniform reference model. Consequently, the term $\beta\left(\log\pi_{\text{ref}}(y_{w}|x)-\log\pi_{\text{ref}}(y_{l}|x)\right)$ collapses to a constant, reflecting the difference in selection probabilities induced by the scoring mechanism. This emphasizes the pivotal role of the implicit reference model in preference optimization.
r  Refer to caption
Figure 1: $\log\pi_{\mathrm{ref}}(y_{w}|x)-\log\pi_{\mathrm{ref}}(y_{l}|x)$ along the training steps. The random fluctuations suggest $\pi_{\mathrm{ref}}$ struggles to distinguish between the preferred and less preferred responses.
Limitations of DPO: As depicted in Figure 1, the reference model $\pi_{\text{ref}}$ in DPO may not effectively distinguish between the preferred ($y_{w}$) and less preferred ($y_{l}$) responses, as its outputs do not inherently reflect the preference information. In contrast, using a uniform distribution as in SimPO results in a reward margin $\gamma$, ensuring that the reward difference between the preferred and less preferred responses is entirely governed by the policy model $\pi_{\theta}$.
Limitations of SimPO: While SimPO simplifies the loss function by using a constant offset $\gamma$, this one-size-fits-all approach overlooks the variability inherent in different data instances. The fixed $\gamma$ across all training samples could lead to suboptimal performance, especially in the presence of noise or inconsistencies in the data. Moreover, completely discarding reference model $\pi_{\text{ref}}$ may eliminate potentially useful prior knowledge about language structure and semantic relationships that could help discriminate between similar response pairs.
2.2 Proposed Method: AlphaDPO
Our analysis highlights the significant impact of the reference model in preference optimization. To overcome the limitations identified in both DPO and SimPO, we propose the following principles:
Principle 1: The reference model should contribute to differentiating between preferred and less preferred responses.
Principle 2: The reference model should adapt to discrepancies between response pairs to capture instance-specific nuances.
Principle 1 addresses the shortcoming in DPO, where the reference model may inadequately distinguish between $y_{w}$ and $y_{l}$, introducing uncertainty without a guaranteed margin. Principle 2 rectifies the oversimplification in SimPO, where the absence of a reference model fails to account for variability across different instances.
Deriving the AlphaDPO Objective. Starting from the standard Reinforcement Learning (RL) objective for preference optimization, we redefine the reference model $\pi_{\text{ref}}$ as an implicit reference model $\hat{\pi}_{\text{ref}}$, formulated as:
$\hat{\pi}_{\text{ref}}(y|x)\propto U(y|x)\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\alpha},\vspace{-15pt}$ (8)
where $\alpha$ is a hyperparameter that determines the extent to which the policy model $\pi_{\theta}$ modifies the reference model $\pi_{\text{ref}}$, and $U(y|x)$ represents a uniform distribution providing a stable baseline. This formulation shares structural similarities with the weak-to-strong preference optimization framework (Zhu et al., 2025), where the aligned strong model $\pi_{r}^{\text{strong}}$ is expressed as:
$\vspace{-15pt}\pi_{r}^{\text{strong}}(y\mid x)\propto\pi_{\text{ref}}^{\text{strong}}(y\mid x)\left(\frac{\pi_{r}^{\text{weak}}(y\mid x)}{\pi_{\text{ref}}^{\text{weak}}(y\mid x)}\right)^{\alpha}.$ (9)
Both formulations employ model ratios to dynamically adjust the reference distribution — $\frac{\pi_{\theta}}{\pi_{\text{ref}}}$ in Equation 8 and $\frac{\pi_{r}^{\text{weak}}}{\pi_{\text{ref}}^{\text{weak}}}$ in Equation 9. Specifically, $\hat{\pi}_{\text{ref}}$ interpolates between existing methods, controlled by the smoothness parameter $\alpha$: i) When $\alpha=0$, $\hat{\pi}_{\text{ref}}$ reduces to the uniform distribution $U(y|x)$, equivalent to the assumption in SimPO. ii) When $\alpha=1$, $\hat{\pi}_{\text{ref}}$ fully incorporates the proportionality $\frac{\pi_{\theta}}{\pi_{\text{ref}}}$, aligning with DPO.
The motivation behind this redefinition is to overcome the potential suboptimality of both SimPO and DPO by learning instance-specific adjustments while preserving useful prior knowledge. Rather than relying on either a fixed reference model or a constant margin, $\hat{\pi}_{\text{ref}}$ creates a flexible framework that can better capture varying degrees of preference strength across different instances. Further details are provided in Appendix C.
Substituting $\hat{\pi}_{\text{ref}}$ into the original DPO loss function, we obtain the AlphaDPO objective:
$\mathcal{L}_{\text{AlphaDPO}}(\pi_{\theta},\pi_{\text{ref}})$ (10)
$=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}-\beta\log\frac{\hat{\pi}_{\text{ref}}(y_{w}|x)}{\hat{\pi}_{\text{ref}}(y_{l}|x)}\right)\right]$
$=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right)-\left[\gamma+\alpha M(x,y_{w},y_{l})\right]\right)\right],$
where $\gamma=\beta\left(\log\frac{U(y_{w}|x)}{U(y_{l}|x)}\right)$ is a constant offset as before, and $M(x,y_{w},y_{l})$ is defined as:
$M(x,y_{w},y_{l})=\beta\left(\log\frac{\pi_{\theta}(y_{w}|x)\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{w}|x)\pi_{\theta}(y_{l}|x)}\right).$ (11)
The term $M(x,y_{w},y_{l})$ measures the divergence between the policy model $\pi_{\theta}$ and the reference model $\pi_{\text{ref}}$ over the response pairs, effectively capturing instance-specific discrepancies as described in Principle 2.
Stop Gradient on $\hat{\pi}_{\text{ref}}$: Although $\hat{\pi}_{\text{ref}}$ depends on $\pi_{\theta}$ and $\pi_{\text{ref}}$, it is intended to serve as a fixed reference during optimization. To prevent gradients from backpropagating through $\hat{\pi}_{\text{ref}}$ to $\pi_{\theta}$, we apply a stop-gradient operation, denoted as $\operatorname{sg}[\cdot]$, ensuring that $\hat{\pi}_{\text{ref}}$ remains constant during the policy updates.
Normalization of $M(x,y_{w},y_{l})$: To stabilize training and avoid $M(x,y_{w},y_{l})$ dominating the loss due to scale variations, we apply Z-score normalization (Patro & Sahu, 2015) to $M$:
$M^{\ast}(x,y_{w},y_{l})=\frac{M(x,y_{w},y_{l})-\mu_{M}}{\sigma_{M}},$ (12)
where $\mu_{M}$ and $\sigma_{M}$ are the mean and standard deviation of $M$ computed over the training dataset.
Length-normalized Reward Formulation: Inspired by SimPO, we incorporate length normalization into our method. This adjustment ensures that rewards are scaled appropriately with respect to the length of the sequences, thereby stabilizing the training process. As demonstrated in our experiments (cf. Appendix D.2), we also confirmed that even without length normalization, our method remains effective and continues to show performance improvements.
Final Objective: Incorporating the above considerations, the final AlphaDPO loss function becomes:
$\mathcal{L}_{\text{AlphaDPO}}(\pi_{\theta},\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (13)
$\left[\log\sigma\left(u(x,y_{w},y_{l})-\operatorname{sg}\left[\gamma+\alpha M^{\ast}(x,y_{w},y_{l})\right]\right)\right],$
where $u(x,y_{w},y_{l})=\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log{\pi_{\theta}(y_{l}|x)}$. This formulation ensures balanced influence between the policy and reference models, aligning with Principles 1 and 2. By incorporating the normalized discrepancy term $M^{\ast}(x,y_{w},y_{l})$, AlphaDPO adaptively adjusts the margin between preferred and less preferred responses based on instance-specific differences, enhancing learning.
3 Theoretical Analysis of AlphaDPO
Balancing alignment performance with response diversity is crucial in recent alignment methods (Zeng et al., 2024; Wang et al., 2024a; Ji et al., 2024a). A popular approach is the Token-Level Direct Preference Optimization (TDPO) method (Zeng et al., 2024), which introduces fine-grained control of the KL divergence at the token level. Given a prompt $x$ and preceding tokens $y^{<t}$, the policy $\pi_{\theta}$ generates the next token $z$ by sampling from $\pi_{\theta}(z|x,y^{<t})$.
By mapping the reward model to a token-level format, the TDPO loss is defined as:
$\mathcal{L}_{\text{TDPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$ (14)
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\delta(x,y_{w},y_{l})\right)\right],$
where the margin term $\delta(x,y_{w},y_{l})$ is defined as:
$\delta(x,y_{w},y_{l})=\beta\mathbb{D}_{\text{SeqKL}}[x,y_{l};\pi_{\text{ref}}||\pi_{\theta}]-\beta\mathbb{D}_{\text{SeqKL}}[x,y_{w};\pi_{\text{ref}}||\pi_{\theta}],$ (15)
$=\beta\sum_{t=1}^{|y_{l}|}\mathbb{E}_{z}[\log\frac{\pi_{\text{ref}}(z|[x,y^{<t}_{l}])}{\pi_{\theta}(z|[x,y^{<t}_{l}])}]-\beta\sum_{t=1}^{|y_{w}|}\mathbb{E}_{z}[\log\frac{\pi_{\text{ref}}(z|[x,y^{<t}_{w}])}{\pi_{\theta}(z|[x,y^{<t}_{w}])}].$ (16)
Here, $z\sim\pi_{\text{ref}}$ and $\mathbb{D}_{\text{SeqKL}}[x,y;\pi_{\text{ref}}||\pi_{\theta}]$ denotes the sequential KL divergence between $\pi_{\text{ref}}$ and $\pi_{\theta}$ along the sequence $y$ given $x$.
Below we present a lemma establishing theoretical connections between TDPO and AlphaDPO w.r.t. the margin terms.
Lemma 3.1 (Equivalence of Margin Terms).
Consider the margin term $\delta(x,y_{w},y_{l})$ that represents the difference in sequential KL divergences between a reference policy $\pi_{ref}$ and policy $\pi_{\theta}$ for preferred sequence $y_{w}$ and rejected sequence $y_{l}$, conditioned on input x:
$\delta(x,y_{w},y_{l})$ $=\beta\mathbb{D}_{\text{SeqKL}}[x,y_{l};\pi_{\text{ref}}||\pi_{\theta}]$ (17)
$\quad-\beta\mathbb{D}_{\text{SeqKL}}[x,y_{w};\pi_{\text{ref}}||\pi_{\theta}],$
Under the assumption that the sequential KL divergence can be approximated by the log-likelihood ratio of complete sequences, the margin term $\delta(x,y_{w},y_{l})$ admits the following approximation:
$\delta(x,y_{w},y_{l})$ $\approx\beta\left(\log\dfrac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\dfrac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
$=M(x,y_{w},y_{l}),$
where $M(x,y_{w},y_{l})$ is the margin term in the AlphaDPO objective.
The proof follows from the sequence-level approximation of the KL divergence between $\pi_{\text{ref}}$ and $\pi_{\theta}$ along a sequence $y$, where:
$\mathbb{D}_{\text{SeqKL}}[x,y;\pi_{\text{ref}}||\pi_{\theta}]$ $=\sum_{t=1}^{|y|}\mathbb{E}_{z\sim\pi_{\text{ref}}}\left[\log\dfrac{\pi_{\text{ref}}(z|x,y^{<t})}{\pi_{\theta}(z|x,y^{<t})}\right]$
$\approx\log\dfrac{\pi_{\text{ref}}(y|x)}{\pi_{\theta}(y|x)}.$
Applying this approximation to both $y_{w}$ and $y_{l}$, the difference $\delta(x,y_{w},y_{l})$ simplifies to the difference of log-probability ratios, thereby establishing the equivalence with the margin term in AlphaDPO.
Lemma 3.1 highlights that the margin term $\delta(x,y_{w},y_{l})$, which represents the sequential KL divergence difference between preferred and rejected responses, can be directly mapped to the term $M(x,y_{w},y_{l})$ in AlphaDPO. This mapping underscores the theoretical connection between the two approaches in terms of alignment control. While TDPO operates at the token level and provides fine-grained control, AlphaDPO offers greater computational efficiency by operating at the sequence level without sacrificing performance (cf. Appendix Table 6). Moreover, the sequence-level approximation enhances robustness to token-level noise in $\pi_{\text{ref}}$, making AlphaDPO particularly suited for scenarios where the reference policy may not be perfectly aligned.
Advantages of the margin term $\delta(x,y_{w},y_{l})$. The core contribution of TDPO lies in introducing the margin term $r(x,y_{w})-r(x,y_{l})-\delta(x,y_{w},y_{l})$, which is similar to DPO with an offset (Amini et al., 2024) and helps control the KL divergence. In contrast, AlphaDPO generalizes this approach by replacing $\delta(x,y_{w},y_{l})$ with $M(x,y_{w},y_{l})$, a more flexible margin term inspired by SimPO. Appendix Table 6 supports this with performance comparisons between TDPO and AlphaDPO. These findings illustrate: i) Adding an offset to DPO and its variants is a robust strategy, applicable to both TDPO and AlphaDPO. ii) The choice of offset is critical — the sequence-level margin term $M(x,y_{w},y_{l})$ is particularly effective when applied to potentially unreliable reference models.
4 Experiments
Table 1: AlpacaEval 2, Arena-Hard results across four settings. “WR” denotes the raw win rate,“LC” the length-controlled win rate, and “SC” the style-controlled win rate. The best results are highlighted in bold, while the second-best are underlined.
Method Llama3-Instruct (8B) Mistral-Instruct (7B)
AlpacaEval 2 Arena-Hard AlpacaEval 2 Arena-Hard
LC (%) WR (%) SC (%) LC (%) WR (%) LC (%) WR (%) SC (%) LC (%) WR (%)
SFT 24.0 23.6 22.1 22.2 22.4 19.0 15.4 18.3 13.2 12.9
DPO 40.2 38.1 31.9 32.0 31.2 20.3 17.9 18.9 13.7 13.4
IPO 35.9 34.4 29.2 29.9 30.2 22.3 18.6 22.4 16.6 16.2
CPO 29.6 34.4 26.3 28.1 29.4 26.2 31.7 26.6 21.4 23.8
KTO 38.3 34.1 30.3 30.6 30.3 19.4 20.3 21.5 16.0 16.8
ORPO 31.6 29.8 26.6 26.6 26.3 24.0 23.0 24.4 18.5 18.6
R-DPO 40.3 37.3 33.1 32.9 32.9 21.4 22.2 18.7 14.0 13.8
SimPO 43.8 38.0 33.5 33.5 32.6 30.2 32.1 25.6 19.8 20.1
AlphaDPO 46.6 38.1 34.1 34.2 33.3 32.3 32.6 27.2 21.5 21.5
Method Llama3-Instruct v0.2 (8B) Gemma2-Instruct (9B)
AlpacaEval 2 Arena-Hard AlpacaEval 2 Arena-Hard
LC (%) WR (%) SC (%) LC (%) WR (%) LC (%) WR (%) SC (%) LC (%) WR (%)
SFT 24.0 23.6 22.1 22.2 22.4 48.7 36.5 32.0 42.2 42.1
DPO 51.9 50.8 26.1 31.5 33.9 70.4 66.9 43.9 55.6 58.8
IPO 40.6 39.6 31.1 34.2 34.9 62.6 58.4 41.1 51.9 53.5
CPO 36.5 40.8 29.4 32.8 34.2 56.4 53.4 42.4 53.3 55.2
KTO 41.4 36.4 27.1 29.5 28.9 61.7 55.5 41.7 52.3 53.8
ORPO 36.5 33.1 28.8 30.8 30.4 56.2 46.7 35.1 45.3 46.2
R-DPO 51.6 50.7 29.2 34.3 35.0 68.3 66.9 45.1 55.9 57.9
SimPO 55.6 49.6 28.5 34.0 33.6 72.4 65.0 45.0 56.1 57.8
AlphaDPO 58.7 51.1 30.8 36.3 35.7 73.4 66.1 48.6 59.3 60.8
In this section, we present the main results of our experiments, highlighting the superior performance of AlphaDPO over existing methods on various benchmarks and ablation studies to analyze the impact of different components of AlphaDPO.
4.1 Experiments Setup
Models and training settings. We optimize preferences using three model families: Llama3-8B (AI@Meta, 2024), Mistral2-7B (Jiang et al., 2023), and Gemma2-9B (Rivière et al., 2024), all in the Instruct setup. Following Meng et al. (2024), we utilize pre-trained instruction-tuned models (meta-llama/Meta-Llama-3-8B-Instruct, mistralai/Mistral-7B-Instruct-v0.2, google/gemma-2-9b-it) as SFT models. For a fair comparison, we use the same training data as SimPO: princeton-nlp/llama3-ultrafeedback-armorm [^1] https://huggingface.co/datasets/princeton-nlp/llama3-ultrafeedback-armorm, princeton-nlp/mistral-instruct-ultrafeedback222 https://huggingface.co/datasets/princeton-nlp/mistral-instruct-ultrafeedback, and princeton-nlp/gemma2-ultrafeedback-armorm  [^3] https://huggingface.co/datasets/princeton-nlp/gemma2-ultrafeedback-armorm for Llama3-8B, Mistral2-7B, and Gemma2-9B, respectively. Additionally, the v0.2 Llama3-Instruct setup uses RLHFlow/ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024b) as the reward model for ranking generated data, significantly enhancing performance. These configurations represent state-of-the-art methods, positioning our models among the top performers on various leaderboards.
Table 2: Ablation studies under Llama3-Instruct v0.2 and Mistral-Instruct settings. We ablate each key design of AlphaDPO and explore variants of the implicit reference model $\hat{\pi}_{\text{ref}}$.
Method Llama3-Instruct v0.2 (8B) Mistral-Instruct (7B)
AlpacaEval 2 Arena-Hard AlpacaEval 2 Arena-Hard
LC (%) WR (%) SC (%) LC (%) WR (%) LC (%) WR (%) SC (%) LC (%) WR (%)
$U(\cdot|x)$ 55.6 49.6 28.5 34.0 33.6 30.2 32.1 25.6 19.8 20.1
$U(\cdot|x)\left({\pi_{\theta}(\cdot|x)}/{\pi_{\text{ref}}(\cdot|x)}\right)^{\alpha}$ 58.7 51.1 30.8 36.3 35.7 32.3 32.6 27.2 21.5 21.5
w/o Normalization 56.5 49.7 23.1 28.4 27.7 32.1 33.1 25.2 19.7 19.6
w/o sg 2.7 3.7 7.7 5.4 6.3 27.2 27.7 25.8 20.3 20.7
$\gamma=0$ 51.2 44.9 30.0 34.5 33.3 31.9 31.3 24.2 19.6 19.3
$U(\cdot|x)\left({\pi_{\theta}(\cdot|x)}\right)^{\alpha}$ 57.2 50.4 27.6 33.5 32.9 31.6 34.1 26.9 21.3 21.5
$U(\cdot|x)\left({\pi_{\text{ref}}(\cdot|x)}\right)^{\alpha}$ 56.3 49.5 29.0 34.3 33.5 28.6 30.9 25.5 20.1 20.3
$U(\cdot|x)\left(1/{\pi_{\text{ref}}(\cdot|x)}\right)^{\alpha}$ 56.3 49.2 29.0 34.4 33.8 32.2 33.1 26.0 20.7 20.6
Evaluation benchmarks. We evaluate our models using two widely recognized open-ended instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023) and Arena-Hard (Li et al., 2024). These benchmarks assess the models’ conversational abilities across a diverse range of queries and are extensively used by the research community. For AlpacaEval 2, we report the length-controlled win rate (LC) and raw win rate (WR). For Arena-Hard, we provide the win rate (WR), length-controlled win rate (LC), and style-controlled win rate (SC) compared to baseline models. Note that style significantly impacts performance on these leaderboards.
Baselines. We compare AlphaDPO with several state-of-the-art preference optimization methods: DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), IPO (Azar et al., 2024), CPO (Xu et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), and R-DPO (Park et al., 2024). We also include the SFT model as a baseline. We thoroughly tune the hyperparameters for each baseline and report the best performance. Further details can be found in Appendix D.1.
4.2 Main Results
AlphaDPO consistently outperforms existing preference optimization methods. As shown in Table 1, while all preference optimization algorithms improve over the SFT baseline, AlphaDPO achieves superior performance compared to existing methods specifically on the AlpacaEval 2 LC metric. These significant improvements highlight the robustness and effectiveness of AlphaDPO. Specifically, AlphaDPO outperforms the best baseline by an average of 3 percentage points in AlpacaEval 2 LC win rate. Furthermore, on benchmarks such as Arena-Hard, AlphaDPO achieves state-of-the-art or second-best results, demonstrating its competitiveness across different evaluation settings.
Impact of Metrics on Leaderboard Rankings. While both benchmarks are widely used, the standard win rate (WR) metric shows poor separability among different methods, making it challenging to distinguish their relative performance. Minor differences in WR may stem from biases towards generating detailed or aesthetically pleasing responses, aligning with observations by Dubois et al. (2024) and Chen et al. (2024a). In contrast, the length-controlled (LC) and style-controlled (SC) win rates offer more reliable and interpretable metrics, as they reduce the influence of verbosity and stylistic biases, thereby better reflecting true performance.
The importance of the design on the implicit reference model. As the core contribution of this work is to propose a novel reference model $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left({\pi_{\theta}(y|x)}/{\pi_{\text{ref}}(y|x)}\right)^{\alpha}$, we also evaluate other variants of the reference model. Specifically, we compare AlphaDPO with three variants: (1) $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left({\pi_{\theta}(y|x)}\right)^{\alpha}$, (2) $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left({\pi_{\text{ref}}(y|x)}\right)^{\alpha}$, and (3) $\hat{\pi}_{\text{ref}}(y|x)=U(y|x)\left(1/{\pi_{\text{ref}}(y|x)}\right)^{\alpha}$. As shown in Table 2, most of the variants perform better than SimPO ($\hat{\pi}_{\text{ref}}(y|x)=U(y|x)$), which demonstrates the importance of adaptive margin between pairs. Besides, our proposed reference model consistently outperforms other variants, indicating the effectiveness of the proposed design.
All key designs in AlphaDPO are crucial. To further analyze the impact of different components in AlphaDPO, we conduct ablation studies by removing key components from AlphaDPO. As shown in Table 2, removing normalization, stop gradient, or setting $\gamma=0$ all lead to significant performance drops, highlighting the importance of these components in AlphaDPO.
4.3 KL divergence control in AlphaDPO
Outstanding Performance and Lower KL. As noted in Rafailov et al. (2023); Zeng et al. (2024), it is crucial to consider both performance and KL divergence when comparing algorithms. A slightly higher win rate accompanied by a significantly higher KL divergence is often not desirable. In line with the design principles of TDPO, we implemented SimPO and AlphaDPO. Figure 2(a) 2(b) presents the KL divergence curves. The results indicate that as $\alpha$ increases, the KL divergence of AlphaDPO remains stable or even decreases slightly when compared to SimPO. This demonstrates that AlphaDPO not only achieves superior performance but also maintains a lower KL divergence, indicating a better balance between alignment and control of KL divergence during the training process.
(a) KL divergence of $y_{w}$ over steps.
(b) KL divergence of $y_{l}$ over steps.
(c) LC and KL divergence.
Figure 2: Analysis of KL divergence and LC trade-off. (a) KL divergence for chosen samples ($y_{w}$), (b) KL divergence for rejected samples ($y_{l}$), and (c) relationship between LC and KL divergence.
Figure 3: Impact of $\alpha$ on (a) LC and SC win rate, (b) reward difference distribution, and (c) log-likelihood distribution of chosen responses in AlphaDPO.
Mitigating Over-Optimization. Over-optimization, as described by Gao et al. (2023) and Rafailov et al. (2024), refers to a phenomenon where model performance exhibits a hump-shaped pattern across different targets: beyond an optimal point, further increasing the KL budget results in diminishing returns. To investigate this, we evaluate SimPO and AlphaDPO at four intermediate checkpoints, corresponding to different KL budgets. As illustrated in Figure 2(c), it is intriguing that while the performance of our approach does decrease with increasing KL budget, the decline is relatively modest. This indicates that our method effectively mitigates the issue of over-optimization.
4.4 The Impact of $\alpha$ in AlphaDPO
Effect of $\alpha$ on Performance. We investigated how the parameter $\alpha$ in AlphaDPO impacts the win rate on AlpacaEval 2 and Arena-Hard. The results, as shown in Figure 3 (a), indicate that the style-control win rate on Arena-Hard initially increases and then decreases with increasing $\alpha$. In contrast, the length-control win rate on AlpacaEval 2 exhibits a consistently increasing trend. This suggests that the optimal value of $\alpha$ varies depending on the evaluation benchmarks. Further experiments refer to Appendix D.3.
Impact of $\alpha$ on the reward distribution. We visualize the distribution of the learned reward margin $r(x,y_{w})-r(x,y_{l})$ and the log likelihood of the chosen response $\log\pi_{\theta}(y_{w}|x)$ under different $\alpha$ values in Figure 3 (b,c). Decreasing $\alpha$ results in a flatter reward margin, while the log likelihood distribution remains relatively unchanged. Conversely, in SimPO (cf. Figure 5), increasing $\gamma$ yields a flatter reward margin distribution but at the cost of also flattening the log likelihood distribution, which undesirably lowers the log likelihood of positive samples. This indicates that AlphaDPO can better balance the relationship between the reward margin and log likelihood.
5 Discussion
Conclusion. We proposed AlphaDPO, an adaptive preference optimization method that improves LLM alignment by introducing a dynamic reward margin based on instance-specific differences. AlphaDPO addresses limitations in previous methods like DPO and SimPO by balancing alignment and diversity through KL divergence control. Our theoretical guarantees and empirical results show that AlphaDPO consistently outperforms baselines on benchmarks like AlpacaEval 2 and Arena-Hard, with significant improvements in win rates, establishing it as a robust solution for LLM fine-tuning.
Limitations and Future Work. While AlphaDPO enhances performance, it introduces an additional hyperparameter, $\alpha$, requiring manual tuning. Future work could focus on developing an adaptive approach to automatically adjust this parameter. Additionally, although we show AlphaDPO’s theoretical equivalence to online methods, it remains an offline approach. Extending it to online learning would allow real-time adaptation, broadening its application in interactive environments. Lastly, we observed that different benchmarks, such as AlpacaEval 2 and Arena-Hard, require distinct parameter settings for optimal performance. Investigating a more generalized approach that adapts effectively across multiple benchmarks would further improve the model’s versatility.
Impact Statement
This work introduces AlphaDPO, an adaptive preference optimization framework that improves Large Language Model (LLM) alignment with human preferences. By establishing a more principled and effective mechanism for preference optimization, AlphaDPO contributes to developing more robustly aligned LLMs, a crucial step towards safer and more beneficial AI systems.
Acknowledgment
This research is supported by the National Science and Technology Major Project (2023ZD0121102), and the National Natural Science Foundation of China (62302321, U24B20180). This research was also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
Appendix A Related Work
Reinforcement learning from human feedback. RLHF is a technique that aligns large language models with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Azar et al., 2024). Traditional RLHF can be divided into three stages: supervised fine-tuning (Zhou et al., 2023; Taori et al., 2023; Geng et al., 2023; Conover et al., 2023; Köpf et al., 2023; Ding et al., 2023), reward modeling (Gao et al., 2023; Luo et al., 2025; Chen et al., 2024b; Lightman et al., 2024; Havrilla et al., 2024; Lambert et al., 2025), and policy optimization (Schulman et al., 2017; Anthony et al., 2017). In the third stage, Proximal Policy Optimization (PPO) is a widely used algorithm. Additionally, Xiong et al. (2023) proposed efficient algorithms for the reverse-KL regularized contextual bandit framework in RLHF. Ye et al. (2024) introduced provably efficient algorithms for KL-regularized Nash-Learning from Human Feedback (NLHF). Furthermore, Ji et al. (2024b) developed an active-query-based PPO algorithm with specific regret bounds and query complexity.
Offline direct preference optimization. Several alternative preference optimization objectives have been proposed in addition to DPO (Rafailov et al., 2023). IPO (Azar et al., 2024) addresses the overfitting issues associated with DPO. ORPO (Hong et al., 2024) and SimPO (Meng et al., 2024) aim to eliminate the dependence on a reference model. R-DPO (Park et al., 2024) focuses on mitigating exploitation based on sequence length. KTO (Ethayarajh et al., 2024) deals with preference optimization when data are not pairwise. CPO (Xu et al., 2024) and $\beta$-DPO(Wu et al., 2024) emphasize the quality of preference data. Another line of research explores comparisons among more than two instances (Dong et al., 2023; Liu et al., 2025a; Song et al., 2024; Yuan et al., 2023).
Online direct preference optimization. Offline direct preference optimization methods are simple but rely on preference data collected offline. RLHF methods interact online with the language model being aligned but require policy gradients. In contrast, online direct preference optimization methods combine the advantages of both approaches. Yuan et al. (2024) proposed a “self-rewarding” approach in which the policy being aligned provides online feedback to itself. Alternatively, OAIF (Guo et al., 2024) is a novel online preference optimization method that can leverage feedback from any LLM, including those stronger than the LLM being aligned. Swamy et al. (2024) also concurrently investigate the importance of online preference but still rely on reward models (RMs). SELMA (Zhang et al., 2025) improves exploration efficiency by selectively favoring responses with high potential rewards rather than indiscriminately sampling unseen responses.
Appendix B Proofs
B.1 Proof of Theorem 2.1
See 2.1
Proof.
Let $U(y|x)$ denote a uniform distribution over the vocabulary $\mathcal{V}$ for a given input $x$. Specifically, for any sequence $y$, the uniform distribution is defined as:
$U(y|x)=\prod_{t=1}^{|y|}\frac{1}{|\mathcal{V}|}=\left(\frac{1}{|\mathcal{V}|}\right)^{|y|}.$
Consider the DPO loss function:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right].$
By substituting $\pi_{\text{ref}}=U$, the term involving the reference policy simplifies to:
$\beta\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}=\beta\left(\log U(y_{w}|x)-\log U(y_{l}|x)\right)=\gamma,$
where $\gamma$ is a constant. This constancy arises because $y_{w}$ and $y_{l}$ are chosen from distinct subsets of the vocabulary, ensuring that $\log U(y_{w}|x)-\log U(y_{l}|x)$ does not depend on the lengths of the sequences but is instead determined by the fixed probabilities of the respective subsets. Consequently, $\gamma$ remains fixed across all samples in $\mathcal{D}$.
Substituting back into the DPO loss function, we obtain:
$\mathcal{L}(\pi_{\theta};U)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\pi_{\theta}(y_{w}|x)-\log\pi_{\theta}(y_{l}|x)\right)-\gamma\right)\right].$
Under the length-normalized reward formulation, the rewards are adjusted by the lengths of the sequences $y_{w}$ and $y_{l}$. This normalization yields:
$\mathcal{L}_{\text{LN}}(\pi_{\theta};U)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right].$
Here, $\gamma$ remains a fixed constant since it is derived from the uniform distribution over distinct vocabulary subsets corresponding to $y_{w}$ and $y_{l}$.
Comparing this with the SimPO loss function:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right],$
it is evident that:
$\mathcal{L}_{\text{LN}}(\pi_{\theta};U)=\mathcal{L}_{\text{SimPO}}(\pi_{\theta}).$
Thus, when the reference policy $\pi_{\text{ref}}$ is a uniform distribution over distinct vocabulary subsets for $y_{w}$ and $y_{l}$, the DPO loss function simplifies to the SimPO loss function with $\gamma$ being a fixed constant. This establishes that SimPO is a special case of DPO under the specified conditions. ∎
B.2 Proof of Lemma 3.1
See 3.1
Proof of Lemma 3.1.
We begin by expanding the definition of $\delta(x,y_{w},y_{l})$:
$\delta(x,y_{w},y_{l})=\beta\mathbb{D}_{\text{SeqKL}}[x,y_{l};\pi_{\text{ref}}\|\pi_{\theta}]-\beta\mathbb{D}_{\text{SeqKL}}[x,y_{w};\pi_{\text{ref}}\|\pi_{\theta}]$
Expanding each sequential KL divergence, we have:
$\mathbb{D}_{\text{SeqKL}}[x,y;\pi_{\text{ref}}\|\pi_{\theta}]=\sum_{t=1}^{|y|}\mathbb{E}_{z\sim\pi_{\text{ref}}}\left[\log\frac{\pi_{\text{ref}}(z\mid[x,y^{<t}])}{\pi_{\theta}(z\mid[x,y^{<t}])}\right]$
Substituting this into the expression for $\delta$, we obtain:
$\delta(x,y_{w},y_{l})=\beta\sum_{t=1}^{|y_{l}|}\mathbb{E}_{z\sim\pi_{\text{ref}}}\left[\log\frac{\pi_{\text{ref}}(z\mid[x,y^{<t}_{l}])}{\pi_{\theta}(z\mid[x,y^{<t}_{l}])}\right]-\beta\sum_{t=1}^{|y_{w}|}\mathbb{E}_{z\sim\pi_{\text{ref}}}\left[\log\frac{\pi_{\text{ref}}(z\mid[x,y^{<t}_{w}])}{\pi_{\theta}(z\mid[x,y^{<t}_{w}])}\right]$
Under the assumption that the reference policy $\pi_{\text{ref}}$ has large errors, we approximate the expectation $\mathbb{E}_{z\sim\pi_{\text{ref}}}$ with a uniform distribution. This approximation simplifies each expectation term as follows:
$\sum_{t=1}^{|y|}\mathbb{E}_{z\sim\pi_{\text{ref}}}\left[\log\frac{\pi_{\text{ref}}(z\mid[x,y^{<t}])}{\pi_{\theta}(z\mid[x,y^{<t}])}\right]\approx\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$
Applying this approximation to both sequential KL divergence terms, we obtain:
$\delta(x,y_{w},y_{l})\approx\beta\left(\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}-\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}\right)$
This expression can be rewritten as:
$\delta(x,y_{w},y_{l})\approx\beta\left(\log\dfrac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\log\dfrac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)=M(x,y_{w},y_{l})$
where $M(x,y_{w},y_{l})$ is the margin term defined in the AlphaDPO objective. Thus, we have shown that:
$\delta(x,y_{w},y_{l})\approx M(x,y_{w},y_{l})$
This completes the proof. ∎
Appendix C The motivation for the proposed $\hat{\pi}_{\text{ref}}(y|x)$
The motivation for the proposed reference policy $\hat{\pi}_{\text{ref}}(y|x)$ can be clarified as follows:
  • Utility Theory Perspective: The proposed $\hat{\pi}_{\text{ref}}(y|x)$ is designed with the uniform distribution $U(y|x)$ as a baseline. The term $\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\alpha}$ dynamically adjusts the reward margin by balancing contributions from the policy and reference models. This mechanism can be interpreted through the lens of utility theory as relative attractiveness, enabling adaptive instance-specific reward modeling.
  • Gradient Perspective By introducing $\hat{\pi}_{\text{ref}}(y|x)$, the framework mitigates the label flipping issues found in DPO or SimPO. In the SimPO framework, the gradient is expressed as:
    $\nabla_{\theta}\mathcal{L}_{\mathrm{SimPO}}(\pi_{\theta})=-\beta\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[s_{\theta}\left(\frac{1}{|y_{w}|}\nabla_{\theta}\log\pi_{\theta}(y_{w}|x)-\frac{1}{|y_{l}|}\nabla_{\theta}\log\pi_{\theta}(y_{l}|x)\right)\right],$
    where $s_{\theta}=\sigma\left(\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)+\gamma\right)$.
    This formulation may amplify weights when the reward estimate is incorrect. By contrast, under AlphaDPO:
    $s_{\theta}=\sigma\left(\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)+\gamma+\alpha M(x,y_{w},y_{l})\right),$
    the additional $\alpha M(x,y_{w},y_{l})$ component increases weight when the reward estimate is accurate, ensuring a more robust reward signal.
  • Motivational Core The central goal of the proposed AlphaDPO is to address the unreliability of the reference policy, as outlined in Section 3.1. By integrating the policy model into the reference model design, the quality of the reference model is enhanced, improving fine-tuning performance. Similar concepts have been explored in recent works (Gorbatovski et al., 2025; Liu et al., 2025b).
Appendix D Experiments
D.1 Implementation Details
We observed that the performance of various methods is highly sensitive to model parameters and learning rates. To ensure a fair comparison, we conducted a hyperparameter search as specified in the respective papers. The specific search ranges are detailed in Table 3. Furthermore, due to recent updates to both Llama3-8b and Instruct-7b models, we had to re-implement SimPO as the original results were no longer directly applicable.
Training hyperparameters.
For other parameters, we used a consistent batch size of 128 across all methods. The learning rate was searched within the range of [3e-7, 5e-7, 8e-7, 1e-6], and all models were trained for a single epoch with a cosine learning rate schedule and a 10% warmup phase. Adam was used as the optimizer (Kingma & Ba, 2014). Additionally, the maximum sequence length was set to 2048.
Table 3: Various preference optimization objectives and hyperparameter search range.
Method Objective Hyperparameter
DPO (Rafailov et al., 2023) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$ $\beta\in[0.01,0.05,0.1]$
IPO (Azar et al., 2024) $\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\tau}\right)^{2}$ $\tau\in[0.01,0.1,0.5,1.0]$
CPO (Xu et al., 2024) $-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)-\lambda\log\pi_{\theta}(y_{w}|x)$ $\alpha=1.0,\,\,\beta\in[0.01,0.05,0.1]$
KTO (Ethayarajh et al., 2024) $-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)+\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$ $\lambda_{l}=\lambda_{w}=1.0$
$\text{where}\,\,z_{\text{ref}}=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\beta\text{KL}\left(\pi_{\theta}(y|x)||\pi_{\text{ref}}(y|x)\right)\right]$ $\beta\in[0.01,0.05,0.1]$
ORPO (Hong et al., 2024) $-\log p_{\theta}(y_{w}|x)-\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right),\,$ $\lambda\in[0.1,0.5,1.0,2.0]$
$\text{where}\,\,p_{\theta}(y|x)=\exp\left(\frac{1}{|y|}\log\pi_{\theta}(y|x)\right)$
R-DPO (Park et al., 2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$ $\alpha\in[0.05,0.1,0.5,1.0]$
$\beta\in[0.01,0.05,0.1]$
SimPO (Meng et al., 2024) $-\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$ $\beta\in[2.0,4.0,6.0,8.0]$
$\gamma\in[0.3,0.5,1.0,1.2,1.4,1.6]$
AlphaDPO $-\log\sigma\left(u(x,y_{w},y_{l})-\operatorname{sg}\left[\gamma+\alpha M^{\ast}(x,y_{w},y_{l})\right]\right)$ $\beta\in[2.5,10.0],\,\,\gamma\in[0.1,0.3,0.5]$
$\text{where}\,\,u(x,y_{w},y_{l})=\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)$ $\alpha\in[1e-2,5e-2,0.1,0.2]$
Table 4: The hyperparameter values in AlphaDPO used for each training setting.
Setting $\beta$ $\gamma$ $\alpha$ Learning rate
Mistral-Instruct 2.5 0.15 5e-2 6e-7
Llama3-Instruct 2.5 0.6 0.2 1e-6
Llama3-Instruct-v0.2 10.0 0.4 0.2 1e-6
Gemma2-Instruct 10.0 0.4 5e-2 8e-7
Hyperparameter in AlphaDPO.
Table 4 outlines the hyperparameters used for AlphaDPO under various settings. It’s important to note that while our approach involves three key parameters, we have found through experience that $\beta$ can be reliably set to 10.0 by default. Among these parameters, $\gamma$ typically requires more careful tuning. As for $\alpha$, we have observed consistent performance improvements when set to 5e-2 by default. If you are already familiar with the parameter settings for SimPO, you can focus your search primarily on $\alpha$ or simply adopt the default setting of $\alpha=5e-2$.
Decoding hyperparameters.
The decoding hyperparameters used in this study are the same as those employed by SimPO [^4] https://github.com/princeton-nlp/SimPO/tree/main/eval. We extend our sincere gratitude to the SimPO team for sharing their invaluable insights.
Computation environment.
All training experiments presented in this paper were conducted using 8×A100 GPUs, as per the procedures detailed in the alignment-handbook repository. [^5] https://github.com/huggingface/alignment-handbook
D.2 AlphaDPO without length-normalized
In this paper, we consider length-normalized training as a stability technique and not as a primary contribution of this work. Existing research (Meng et al., 2024) has demonstrated that length normalization can indeed enhance model performance, particularly with respect to the length control win rate. However, to validate the general applicability of AlphaDPO—specifically, its stability and performance without length normalization—we conducted experiments across several models: meta-llama/Meta-Llama-3-8B-Instruct, mistralai/Mistral-7B-Instruct-v0.2, and google/gemma-2-9b-it.
Figure 4: AlphaDPO LC on AlpacaEval 2 with different $\alpha$ values.
We evaluated DPO, SimPO without length normalization, and AlphaDPO without length normalization. The experimental results, as shown in Table 5, demonstrate that AlphaDPO consistently achieves performance improvements even without the use of length normalization. This indicates the robustness and general effectiveness of AlphaDPO.
Table 5: Performance comparison without length-normalization on AlpacaEval2. “LC” denotes the length-controlled win rate, and “WR” represents the raw win rate.
Method Llama3-Instruct (8B) Mistral-Instruct (7B) Llama3-Instruct v0.2 (8B) Gemma2-Instruct (9B)
LC (%) WR (%) LC (%) WR (%) LC (%) WR (%) LC (%) WR (%)
DPO 40.2 38.1 20.3 18.0 51.1 53.3 70.2 66.9
SimPO w/o LN 42.4 40.4 30.5 38.2 49.2 52.6 71.2 69.9
AlphaDPO w/o LN 44.4 42.6 32.0 38.4 51.1 54.0 72.7 70.5
D.3 AlphaDPO with differenct $\alpha$
To analyze the impact of $\alpha$ on the model, we adjust its value for four different models. The results are illustrated in Figure 4. When $\alpha$ is set to 0, the model degenerates to SimPO. As $\alpha$ increases, performance improves across all models, although the optimal value of $\alpha$ varies among them. This highlights the significance of $\alpha$.
It is noteworthy that within the parameter tuning range [1e-2, 5e-2, 0.1, 0.2], the optimal $\alpha$ values are consistently around 0.1 or even closer to 5e-2.
D.4 Comparison With TDPO
To investigate the relationship between TDPO and AlphaDPO, we conducted the experiments, with the results outlined below.
Table 6: Performance comparison between TDPO and AlphaDPO.
Method Llama3-Instruct (8B)
LC (%) WR (%)
TDPO 52.8 45.9
AlphaDPO w/ $\delta(x,y_{w},y_{l})$ 56.9 50.4
AlphaDPO w/ ${M}(x,y_{w},y_{l})$ 58.7 51.1
In its original form, TDPO did not perform well on Llama3-8B. By applying Lemma 3.1, we modified the expression $M(x,y_{w},y_{l})$ in AlphaDPO to use TDPO’s $\delta(x,y_{w},y_{l})$, converting our sentence-level estimations to a token-level calculation. This adjustment resulted in a noticeable performance improvement, which we attribute to the length-normalization, $\gamma$ and z-score normalization of $\delta(x,y_{w},y_{l})$. Nevertheless, the modified TDPO still underperformed compared to AlphaDPO. This indicates that, when the $\pi_{\text{ref}}$ is suboptimal, token-level calculations are prone to significant errors.
Figure 5: (a) SimPO: Reward difference distribution under different $\gamma$ values. (b) AlphaDPO: Reward difference distribution under different $\alpha$ values. (c) SimPO: Log likelihood distribution on chosen responses under different $\gamma$ values. (d) AlphaDPO: Log likelihood distribution on chosen responses under different $\alpha$ values.